# Inertial Confinement Fusion Forecasting via LLMs

### v2

@@ Abstract
Controlled fusion energy is deemed pivotal for the advancement of human civilization. In this study, we introduce Fusion-LLM, a novel integration of Large Language Models (LLMs) with classical reservoir computing paradigms tailored to address challenges in Inertial Confinement Fusion (ICF). Our approach offers several key contributions: Firstly, we propose the LLM-anchored Reservoir, augmented with a fusion-specific prompt, enabling accurate forecasting of hot electron dynamics during implosion. Secondly, we develop Signal-Digesting Channels to temporally and spatially describe the laser intensity across time, capturing the unique characteristics of ICF inputs. Lastly, we design the Confidence Scanner to quantify the confidence level in forecasting, providing valuable insights for domain experts to design the ICF process. Extensive experiments demonstrate the superior performance of our method, achieving 1.90 CAE, 0.14 top-1 MAE, and 0.11 top-5 MAE in predicting Hard X-ray (HXR) energies of ICF tasks, which presents state-of-the-art comparisons against concurrent best systems. Additionally, we present Fusion4AI, the first ICF benchmark based on physical experiments, aimed at fostering novel ideas in plasma physics research and enhancing the utility of LLMs in scientific exploration. Overall, our work strives to forge an innovative synergy between AI and plasma science for advancing fusion energy.
@@ 1 Introduction
-- The Kardashev scale [34]
After National Ignition Facility (NIF) achieving ignition in December 2022 [1], the focus of current inertial confinement fusion (ICF) research shifts to exploring high gain schemes required to make fusion a practical and sustainable energy source for humankind. Fusion represents a potential key enabler for advancing humanity towards a Type I civilization on the Kardashev scale [91], offering a virtually limitless and clean energy source that could power our civilization globally. This advancement could potentially resolve numerous crises we currently face — $e.g.,$ economic recessions and climate change — by eliminating the need for finite resources like fossil fuels.
The optimization of ICF designs and the achievement of reliable ignition face formidable constraints [8, 14] due to laser-plasma instabilities (LPI)[22, 54]. Efficient and symmetrical driving of the target, vital for ICF, is impeded by LPI phenomena such as stimulated Raman and Brillouin backscatterings (SRS and SBS), which can disrupt implosion symmetry and reduce efficiency through cross-beam energy transfer (CBET) [68, 21]. Hot electrons, a byproduct of LPI processes like SRS and Two-Plasmon-Decay (TPD), can both hinder and assist ignition, showcasing the complexity of the interaction. Despite efforts to measure and simulate hot electrons, obtaining predictive scaling laws remains challenging due to the dynamic nature of laser/plasma conditions and computational limitations [38, 57, 89, 66, 42], highlighting the current gap in establishing a predictive capability based on first principles that aligns with experimental data. These constraints highlight the need for an innovative approach to overcome these obstacles.
Figure 1: The overview of Fusion-LLM.
Currently, Large Language Models (LLMs) exhibit versatile capabilities across diverse disciplines ($e.g.,$ robotics [44, 73, 90, 26], medical health [41, 67, 77, 50], agriculture [56, 80]), adeptly capturing intricate patterns in multimodal data. Due to their success in other domains [39, 93, 46], we are convinced that LLMs may also potentially excel in generalizing to plasma physics, particularly in forecasting the behavior of hot electrons generated during implosion in the real-world, physics experiments. Leveraging their vast pre-trained knowledge base, LLMs could optimize ICF designs by efficiently evaluating numerous scenarios, aiding researchers in identifying promising approaches expediently, and generating insights to enhance our understanding and control of the fusion process.
@@
We conceptualize LLMs as a computational reservoir to unlock their potential for robust domain adaptation and generalization for ICF task, titled Fusion-LLM (see Fig. 1). In order to address question , we propose the LLM-anchored Reservoir, augmented with a fusion-specific prompt, to facilitate the interpretation of plasma physics. The incorporated prompts encompass domain-specific knowledge, instructional cues, and statistical information, thereby enabling the LLMs to accommodate the specific demands of the given task. Additionally, we develop the Signal-Digesting Channels to capture the distinctive characteristics of ICF inputs. It features a temporal encoder to better align the laser signals with the pre-trained, time-series space, and a spatial encoder to provide a global description of the input landscape. To tackle question , we introduce the Confidence Scanner to assess the trustworthiness in predictions. Specifically, we couple the gradient saliency in prediction head and token entropy in LLM outputs to obtain the model’s confidence scores.
@@
  • We represent a pioneer investigation into the use of LLMs for analyzing hot electron dynamics in ICF. LLMs offer a cost-effective alternative compared to both plasma-physics experiments and simulations at the ignition scale. Empirical evidence (see §3/S2) showcases the efficacy of the application of LLM for plasma physics study, which directly benefits the practical design of ICF.
  • We construct an LLM-anchored reservoir computing framework to predict the hot electron dynamics in ICF implosions. Compared to prior arts in reservoir computing [43, 20] and time-series-based LLM [31], our approach requires less data (see Table 2d) and shorter training schedule (see Table 2e), while achieving superior performance.
  • We develop the first fusion benchmarkFusion4AI (see §3.1) based on physical experiments, to facilitate new ideas in plasma physics research and the use of LLMs in scientific exploration.
@@ 2.1 Preliminary
The ICF Overview. A ICF shot is depicted in Fig. 2. Laser beams [Uncaptioned image] are directed towards a target fuel pellet [Uncaptioned image], rapidly heating its surface to generate a plasma envelope [Uncaptioned image]. This plasma envelope subsequently undergoes implosion through a rocket-like expulsion of material, compressing the pellet to densities exceeding that of lead by twenty-fold and attaining ignition temperatures of approximately 100,000, [^0] C, thus initiating fusion reactions. Throughout this process, laser-induced plasma instabilities give rise to hot electrons, which may prematurely heat the pellet and emit Hard X-Rays (HXR) emissions [Uncaptioned image] as they interact within the plasma environment. Understanding the hot electron dynamics is crucial for guiding the design of the ICF. In the context, HXR diagnostics have been extensively applied in studies at facilities such as OMEGA [69, 88, 71, 87] and the NIF [58, 59, 70]. It is however worth noting that conducting these experiments is extremely expensive (around \$1m for one shot).
Experiment and Data Collection. To collect the fusion data, we conduct 100 real-world shots of ICF. Each shot is accompanied by detailed configurations including the size of the fuel pellet, the phase plate of the laser, and the input laser profile. We also measure raw sensor readings regarding HXR, which can be further converted into charges of hot electrons, allowing for the subsequent calculation of the total energy of these electrons based on their temperature. In this study, we focus on forecasting the energy from HXR emissions by given laser intensity inputs throughout the shot. This data is indispensable for physicists to ascertain and analyze the state of ICF.
2.2 Fusion-LLM
@@
In this section, we introduce Fusion-LLM, a novel approach for predicting LPI and hot-electron generation in ICF. Illustrated in Fig. 3, the inputs comprise fusion-specific prompts and the time series of input laser intensity, which are processed through LLM for feature extraction. Subsequently, the output module makes predictions of hot-electron energy along with confidence scores. Our approach is characterized by three core modules: LLM-anchored Reservoir, which establishes a reservoir foundation using LLMs to comprehend the dynamic impact of laser intensity on hot-electron energy emission; Signal-Digesting Channels is responsible for encoding the time series of input laser intensity, capturing the temporal characteristics of sequential details, and spatial distinctiveness of data landscapes; and Confidence Scanner, tasked with estimating prediction confidence for each shot, thereby providing trustworthy guidance for practical experimental design of the ICF.
@@
Figure 3: Fusion-LLM framework. (a) Fusion-specific prompts structure domain textual prompts with context, task and input descriptions (see §2.2.1 for details). (b) Signal-digesting channels comprise a pre-trained temporal encoder to extract time-series features of laser signals, and a spatial encoder to encode critical landscapes of the inputs (see §2.2.2 for details). For simplicity, we skip some architectural modules. We provide implementation details in appendix (see §S1).
@@ 2.2.1 LLM-anchored Reservoir
The ${}^{**}\texttt{Prompt}$ specifically denotes the structured FSP, comprising three descriptors (see Fig. 3a) designed with particular focus: ➤ Context descriptor provides a detailed overview of the ICF process, highlighting the nature, sources, and characteristics of the data. It elaborates on the experimental procedures (see §2.1) used in data collection, emphasizing principles and methodologies specific to ICF. This descriptor enhances the LLMs’ comprehension of the experimental context.  ➤ Task descriptor outlines explicit instructions for the prediction tasks, including the format and expected length of the output. It guides the inference and forecasting process by specifying imperative considerations, ensuring that the forecasting align with domain-specific insights. ➤ Input descriptor presents a concise statistical summary of the input data, offering insights into its distribution and key characteristics such as minimum, maximum, and median values. This descriptor is vital for informing the LLMs about the underlying statistical properties of the input signals. Collectively, this prompting strategy facilitates the LLMs’ ability to mine intricate input signals and produce predictions that are scientifically robust and contextually coherent within the phase space of the reservoir.
@@
  • Adaptability for science. The utilization of LLMs in ICF forcasting exhibits notable adaptability in confronting the pivotal scientific challenge – modeling LPI and hot electron generation. Later, we will present empirical evidence to substantiate the systemic efficacy (see §3.1). The success of this methodology is poised to provide a generic alternative for other adjacent scientific domains in pursuit of LLM-based solutions.
@@ 2.2.2 Signal-Digesting Channels
With the establishment of the LLM-based approach described in §2.2.1, robust predictions can be immediately attained (see Table 2c). Due to the uniqueness of the ICF data, the input HXR signal exhibits its time-series, temporal landscape. Recognizing the sensitivity of LLMs to input data [64, 65], we introduce Signal-Digesting Channels (SDC), conceptually aligned with our reservoir design, to capture crucial input characteristics and further augment the performance of LLMs.
@@
Temporal encoder is designed to extract time-series features using a windowing mechanism that constructs consistent signal patches across sequential time steps. It employs a set of Transformer layers [85] to capture time-series distributions over the forecast horizon. This process is formulated as $f_{tmp}:(X_{t-l:t+h},Y_{t-l:t})\mapsto\hat{\psi}$, where $X$ and $Y$ represent the input data and target data spanning time windows of length $l$ and $h$ at time $t$. The encoder is pre-trained on a large-scale time-series dataset (see §S1 for details) and is optimized using the log-likelihood of the forecast:
@@ 2.2.3 Confidence Scanner
Trustworthiness is pivotal for AI in science. Under-confident predictions may lead to misguided conclusions or improper actions. Some approaches [45, 27] directly utilize the entropy observed in the output tokens of the LLMs to gauge the confidence. However, these methods encounter a challenge in our study of ICF whereby the entropy of LLM’s output tokens does not consistently reflect confidence of prediction at each time step. This discrepancy arises due to the non-linear transformation undertaken by the multi-layer perception within the prediction head, which alters the embedding of token counterparts, thereby distorting the relation between tokens and their corresponding confidence estimations.
To this end, we propose Confidence Scanner that incorporates a confidence reweighing mechanism to assess the confidence level of each prediction systematically. Concretely, our approach re-calibrates the allocation of confidence across tokens to implicitly reflect their actual influence on predictions. As shown in Fig. 4, we extract the embedding $E_{k}=\{e_{n-k},\dots,e_{n}\}$ for the last layer of LLM, which specifically analyze the embedding of the last $k$ tokens. The confidence level $H$ is then formulated as:
@@ 3 Empirical Findings
Figure 5: Qualitative results of hot electron prediction. We plot Ground Truth and the predictions of Ours, Time-LLM, LSTM and Autoformer. Y and X axes denote energy and time steps respectively.
@@ 3.1 Main Results
Dataset. We develop a new benchmark, Fusion4AI, to support AI research in ICF. This benchmark consists of 40,000 LPI samples containing 100 shot sequences with 400 time steps/shot. Each shot is documented with key parameters such as target size, laser intensity, and energy of hot electrons (see §2.1). The dataset has been systematically divided into 80/10/10 for train/val/test splits respectively. We will release this dataset upon acceptance to advance research for the fusion reaction.
Baselines. We choose a classic physical Particle-In-Cell (PIC) simulation method [13], a number of classic AI models (i.e., LSTM [25], Autoformer [86]), reservoir computing models (i.e., HoGRC [43], RCRK [17], NGRC [20]), and concurrent time-series LLM-based models (i.e., GPT4TS [94] and Time-LLM [31]) as baseline models for performance comparison on proposed Fusion4AI dataset.
Experimental Setup. Our experiments are trained with 100 epochs and a batch size of 5, which is adequate to achieve convergence based on our empirical findings. In addition, we utilize a fixed learning rate of $0.0004$ and the Adam optimizer [36]. A loss function defined by the cumulative absolute error across each time-steps $f_{\text{loss}}(Y,G)=\sum_{n=1}^{\text{pred\_len}}|y_{n}-g_{n}|$ is used, where $Y$ and $G$ represent the sequences of predictions and ground truths, respectively, $y_{n}$ and $g_{n}$ denote the values at the $n$-th time-step, and pred_len is the length of prediction. For all other counterpart methods, we follow the original experimental setting and training configurations to reproduce the results.
@@
Table 1: Quantitative results on Fusion4AI test split for hot electron energy predictions (see §3.1 for details). Refer to the metrics section for details of CAE, top-1 and top-5 MAE.
@@
PIC Simulation [13] 2.88 0.20 0.13
LSTM [25] 5.82±0.06 0.35±0.01 0.35±0.01
Autoformer [86] 5.79±0.04 0.35±0.01 0.34±0.01
HoGRC [43] 4.20±0.79 0.25±0.05 0.22±0.02
RCRK [17] 4.31±0.46 0.28±0.04 0.22±0.01
NGRC [20] 4.28±0.68 0.27±0.04 0.23±0.02
GPT4TS [94] 3.34±0.58 0.18±0.05 0.14±0.04
Time-LLM [31] 3.48±0.72 0.18±0.05 0.15±0.05
Fusion-LLM (Llama-2-7B) 2.15±0.26 0.14±0.01 0.12±0.01
Fusion-LLM (Llama-3-8B) 1.90±0.33 0.14±0.01 0.11±0.01
Performance Comparison. The primary challenge in forecasting for ICF lies in developing a robust yet efficient predictive model. Table 1 empirically demonstrates the effectiveness of our system in predicting output energy in ICF, showcasing a significant advantage over all other baseline methods across all metrics. Specifically, our approach with Llama-3 surpasses PIC simulation model [13] by 0.98 in terms of CAE. Additionally, compared to classic AI methods, our approach outperforms LSTM [25] and Autoformer [86] by 3.92 and 3.89, respectively. Our margins in CAE over three other reservoir computing methods, HoGRC [43], RCRK [17], NGRC [20], are 2.30, 2.41, and 2.38, respectively. Moreover, our method outperforms concurrent LLM-based approaches GPT4TS [94] and Time-LLM [31] by 1.44 and 1.58 on CAE, ${}_{\!}$respectively, ${}_{\!}$while ${}_{\!}$maintaining ${}_{\!}$comparable ${}_{\!}$speed. ${}_{\!}$etailed ${}_{\!}$analysis ${}_{\!}$is ${}_{\!}$supplemented ${}_{\!}$in ${}_{\!}$§S2.
@@
Figure 6: Visualization of confidence score and prediction error. Samples correspond to Fig. 5.
We further present qualitative results in Fig. 5, aligning with our quantitative findings that our method surpasses all comparative baselines in predictive accuracy. Additionally, Fig. 6 illustrates the confidence scores associated with our predictions. The visualization elucidates a clear correlation between predictive error and confidence scores, indicating high confidence corresponding to low errors and conversely. Notably, our approach consistently demonstrates a heightened level of confidence, particularly in forecasting peak values across the temporal sequence, a critical phase in ICF.
Table 2: A set of ablative studies. The experiments are evaluted on val split.

### v3

# Inertial Confinement Fusion Forecasting via Large Language Models

@@ Abstract
Controlled fusion energy is deemed pivotal for the advancement of human civilization. In this study, we introduce LPI-LLM, a novel integration of Large Language Models (LLMs) with classical reservoir computing paradigms tailored to address a critical challenge, Laser-Plasma Instabilities (LPI), in Inertial Confinement Fusion (ICF). Our approach offers several key contributions: Firstly, we propose the LLM-anchored Reservoir, augmented with a Fusion-specific Prompt, enabling accurate forecasting of LPI-generated-hot electron dynamics during implosion. Secondly, we develop Signal-Digesting Channels to temporally and spatially describe the driver laser intensity across time, capturing the unique characteristics of ICF inputs. Lastly, we design the Confidence Scanner to quantify the confidence level in forecasting, providing valuable insights for domain experts to design the ICF process. Extensive experiments demonstrate the superior performance of our method, achieving 1.90 CAE, 0.14 top-1 MAE, and 0.11 top-5 MAE in predicting Hard X-ray (HXR) energies emitted by the hot electrons in ICF implosions, which presents state-of-the-art comparisons against concurrent best systems. Additionally, we present LPI4AI, the first LPI benchmark based on physical experiments, aimed at fostering novel ideas in LPI research and enhancing the utility of LLMs in scientific exploration. Overall, our work strives to forge an innovative synergy between AI and ICF for advancing fusion energy.
@@ 1 Introduction
-- The Kardashev scale (Kardashev, 1964)
After National Ignition Facility (NIF) achieving ignition in December 2022 (Abu-Shawareb et al., 2024), the focus of current inertial confinement fusion (ICF) research shifts to exploring high gain schemes required to make fusion a practical and sustainable energy source for humankind. Fusion represents a potential key enabler for advancing humanity towards a Type I civilization on the Kardashev scale (Zhang et al., 2023), offering a virtually limitless and clean energy source that could power our civilization globally. This advancement could potentially resolve numerous crises we currently face — $e.g.,$ economic recessions and climate change — by eliminating the need for finite resources like fossil fuels.
Direct-drive ICF has potentially higher gains due to more efficient driver-target coupling but faces many challenges (Betti & Hurricane, 2016). The optimization of ICF designs to achieve reliable high-gain ignition faces formidable constraints (Betti & Hurricane, 2016; Craxton et al., 2015) due to laser-plasma instabilities (LPI(Gopalaswamy et al., 2024; Radha et al., 2016). Efficient and symmetrical driving of the target, vital for ICF, is impeded by LPI phenomena such as stimulated Raman and Brillouin backscatterings (SRS and SBS), which can disrupt implosion symmetry and reduce efficiency through cross-beam energy transfer (CBET) (Smalyuk et al., 2008a; Goncharov et al., 2008). Hot electrons, a byproduct of LPI processes like SRS and Two-Plasmon-Decay (TPD), can both hinder (Smalyuk et al., 2008a; Goncharov et al., 2008; Craxton et al., 2015; Radha et al., 2016) and assist (Betti et al., 2007; Perkins et al., 2009; Shang et al., 2017) ignition, showcasing the importance of LPI. Despite efforts to measure and simulate hot electron generation, obtaining predictive scaling laws remains challenging due to the dynamic nature of laser/plasma conditions and computational limitations (Klimo et al., 2010; Riconda et al., 2011; Yan et al., 2014; Shang et al., 2017; Li et al., 2020), highlighting the current gap in establishing a predictive capability based on first principles that aligns with experimental data. These constraints highlight the need for an innovative approach to overcome these obstacles.
Figure 1: The overview of LPI-LLM.
Currently, Large Language Models (LLMs) exhibit versatile capabilities across diverse disciplines ($e.g.,$ robotics (Lin et al., 2022; Szot et al., 2021; Yao et al., 2023; Huang et al., 2022), medical health (Li et al., 2023; Singhal et al., 2023; Thirunavukarasu et al., 2023; Moor et al., 2023), agriculture (Rezayi et al., 2022; Tzachor et al., 2023)), adeptly capturing intricate patterns in multimodal data. Due to their success in other domains (Kojima et al., 2022; Zheng et al., 2023; Liu et al., 2023), we are convinced that LLMs may also potentially excel in generalizing to plasma physics, particularly in forecasting the behavior of hot electrons generated during implosion in the real-world, physics experiments. Leveraging their vast pre-trained knowledge base, LLMs could optimize ICF designs by efficiently evaluating numerous scenarios, aiding researchers in identifying promising approaches expediently, and generating insights to enhance our understanding and control of the fusion process.
@@
We conceptualize LLMs as a computational reservoir to unlock their potential for robust domain adaptation and generalization for ICF task, titled LPI-LLM (see Fig. 1). In order to address question , we propose the LLM-anchored Reservoir, augmented with a fusion-specific prompt, to facilitate the interpretation of plasma physics. The incorporated prompts encompass domain-specific knowledge, instructional cues, and statistical information, thereby enabling the LLMs to accommodate the specific demands of the given task. Additionally, we develop the Signal-Digesting Channels to capture the distinctive characteristics of ICF inputs. It features a temporal encoder to better align the laser signals with the pre-trained, time-series space, and a spatial encoder to provide a global description of the input landscape. To tackle question , we introduce the Confidence Scanner to assess the trustworthiness in predictions. Specifically, we couple the gradient saliency in prediction head and token entropy in LLM outputs to obtain the model’s confidence scores.
@@
  • We represent a pioneer investigation into the use of LLMs for analyzing hot electron dynamics in ICF. LLMs offer a cost-effective alternative compared to both ICF experiments and plasma-physics simulations. Empirical evidence (see §3/S2) showcases the efficacy of the application of LLM for LPI study, which directly benefits the practical design of ICF.
  • We construct an LLM-anchored reservoir computing framework to predict the hot electron dynamics in ICF implosions. Compared to prior arts in reservoir computing (Li et al., 2024; Gauthier et al., 2021) and time-series-based LLM (Jin et al., 2023), our approach requires less data (see Table 2d) and shorter training schedule (see Table 2e), while achieving superior performance.
  • We develop the first LPI benchmarkLPI4AI (see §3.1) based on physical experiments, to facilitate new ideas in LPI research and the use of LLMs in scientific exploration.
@@ 2.1 Preliminary
The ICF Overview. ICF is one of the two major branches of fusion energy research. The main idea of ICF is using intense drivers to compress the target and maintain the fusion fuel at fusion densities and temperatures with its own inertia, which facilitates the occurrence of fusion reactions. The process of a conventional direct drive ICF is depicted in Fig. 2. Laser beams [Uncaptioned image] are directed towards a target filled with fuel [Uncaptioned image], rapidly heating its surface to create a plasma envelope [Uncaptioned image]. The target surface will eventually blow off, causing the target shell to implode toward the center via the rocket effect. The kinetic energy of the shell will be converted into thermal energy of the fuel, which sustains the fuel at high densities (more than twenty times that of lead) and temperatures (approximately 100,000, [^0] C), thus initiating fusion reactions. Throughout this process, laser plasma instabilities (LPIs) could take energy from the laser and generate hot electrons, which may prematurely heat the target and emit Hard X-Rays (HXR) via bremsstrahlung radiation [Uncaptioned image] as they interact with the surrounding plasma. HXR diagnostics have been extensively applied in experiments at OMEGA (Smalyuk et al., 2008b; Yaakobi et al., 2000; Stoeckl et al., 2003; Yaakobi et al., 2005) and NIF (Rosenberg et al., 2018; 2020; Solodov et al., 2020) to determine the hot electron energy and temperature. It is worth noting that conducting these experiments is expensive (around \$1m for one successful NIF shot). Having a predictive tool will not only improve our understanding of hot electron physics, but also benefit the design of ICF.
Experiment and Data Collection. The HXR data were collected from 100 shots conducted on the OMEGA platform. Detailed configurations, including the target size, the phase plate shape for laser smoothing, and the input laser profile, were available. HXR signals were recorded from four diagnostic channels, enabling the calculation of the total energy and temperature of the electrons. In this study, we focus on forecasting HXR signals with given laser intensity profiles throughout the shot. This data is indispensable for physicists to design new ICF experiments.
2.2 LPI-LLM
@@
In this section, we introduce LPI-LLM, a novel approach for predicting LPI and hot-electron generation in ICF. Illustrated in Fig. 3, the inputs comprise fusion-specific prompts and the time series of input laser intensity, which are processed through LLM for feature extraction. Subsequently, the output module makes predictions of hot-electron energy along with confidence scores. Our approach is characterized by three core modules: LLM-anchored Reservoir, which establishes a reservoir foundation using LLMs to comprehend the dynamic impact of laser intensity on hot-electron energy emission; Signal-Digesting Channels is responsible for encoding the time series of input laser intensity, capturing the temporal characteristics of sequential details, and spatial distinctiveness of data landscapes; and Confidence Scanner, tasked with estimating prediction confidence for each shot, thereby providing trustworthy guidance for practical experimental design of the ICF.
@@
Figure 3: LPI-LLM framework. (a) Fusion-specific prompts structure domain textual prompts with context, task and input descriptions (see §2.2.1 for details). (b) Signal-digesting channels comprise a pre-trained temporal encoder to extract time-series features of laser signals, and a spatial encoder to encode critical landscapes of the inputs (see §2.2.2 for details). For simplicity, we skip some architectural modules. We provide implementation details in appendix (see §S1).
@@ 2.2.1 LLM-anchored Reservoir
The ${}^{**}\texttt{Prompt}$ specifically denotes the structured FSP, comprising three descriptors (see Fig. 3a) designed with particular focus: ➤ Context descriptor provides a detailed overview of the ICF process, highlighting the nature, sources, and characteristics of the data. It elaborates on the experimental procedures (see §2.1) used in data collection, emphasizing principles and methodologies specific to ICF. This descriptor enhances the LLMs’ comprehension of the experimental context.  ➤ Task descriptor outlines explicit instructions for the prediction tasks, including the format and expected length of the output. It guides the inference and forecasting process by specifying imperative considerations, ensuring that the forecasting aligns with domain-specific insights. ➤ Input descriptor presents a concise statistical summary of the input data, offering insights into its distribution and key characteristics such as minimum, maximum, and median values. This descriptor is vital for informing the LLMs about the underlying statistical properties of the input signals. Collectively, this prompting strategy facilitates the LLMs’ ability to mine intricate input signals and produce predictions that are scientifically robust and contextually coherent within the phase space of the reservoir.
@@
  • Adaptability for fusion. The utilization of LLMs in ICF forcasting exhibits notable adaptability in confronting the pivotal scientific challenge – modeling LPI and hot electron generation. Later, we will present empirical evidence to substantiate the systemic efficacy (see §3.1). The success of this methodology is poised to provide a generic alternative for other adjacent scientific domains in pursuit of LLM-based solutions.
@@ 2.2.2 Signal-Digesting Channels
With the establishment of the LLM-based approach described in §2.2.1, robust predictions can be immediately attained (see Table 2c). Due to the uniqueness of the ICF data, the input HXR signal exhibits its time-series, temporal landscape. Recognizing the sensitivity of LLMs to input data (Sclar et al., 2023; 2024), we introduce Signal-Digesting Channels (SDC), conceptually aligned with our design, to capture crucial input characteristics and further augment the performance of LLMs.
@@
Temporal encoder is designed to extract time-series features using a windowing mechanism that constructs consistent signal patches across sequential time steps. It employs a set of Transformer layers (Woo et al., 2024) to capture time-series distributions over the forecast horizon. This process is formulated as $f_{tmp}:(X_{t-l:t+h},Y_{t-l:t})\mapsto\hat{\psi}$, where $X$ and $Y$ represent the input data and target data spanning time windows of length $l$ and $h$ at time $t$. The encoder is pre-trained on a large-scale time-series dataset (see §S1 for details) and is optimized using the log-likelihood of the forecast:
@@ 2.2.3 Confidence Scanner
Trustworthiness is pivotal for AI in science. Under-confident predictions may lead to misguided conclusions or improper actions. Some approaches (Lin et al., 2023; Huang et al., 2023) directly utilize the entropy observed in the output tokens of the LLMs to gauge the confidence. However, these methods encounter a challenge in our study of ICF whereby the entropy of LLM’s output tokens does not consistently reflect confidence of prediction at each time step. This discrepancy arises due to the non-linear transformation undertaken by the multi-layer perception within the prediction head, which alters the embedding of token counterparts, thereby distorting the relation between tokens and their corresponding confidence estimations.
To this end, we propose Confidence Scanner that incorporates a confidence reweighing mechanism to assess the confidence level of each prediction systematically. Concretely, our approach redistributes confidence across tokens to implicitly reflect their actual influence on predictions.
As shown in Fig. 4, we extract the embedding $E_{k}=\{e_{n-k},\dots,e_{n}\}$ for the last layer of LLM, which specifically analyze the embedding of the last $k$ tokens. The confidence level $H$ is then formulated as:
@@ 3 Empirical Findings
Figure 5: Qualitative results of 2 hot electron prediction cases. We plot Ground Truth and the predictions of Ours, Autoformer, Time-LLM and LSTM. Y and X axes denote hot electron energy in voltage, and time steps with step length of 0.025 nanosecond respectively.
@@ 3.1 Main Results
Dataset. We develop a new benchmark, LPI4AI, to support AI research in ICF. This benchmark consists of 40,000 LPI samples containing 100 shot sequences with 400 time steps/shot. Each shot is documented with key parameters such as target size, laser intensity, and energy of hot electrons (see §2.1). The dataset has been systematically divided into 80/10/10 for train/val/test splits respectively. We will release this dataset upon acceptance to advance research for the fusion reaction.
Baselines. We choose a classic physics-based Particle-In-Cell (PIC) simulation method (Cao et al., 2022), a number of classic AI models (i.e., LSTM (Hochreiter & Schmidhuber, 1997), Autoformer (Wu et al., 2021)), reservoir computing models (i.e., HoGRC (Li et al., 2024), RCRK (Dong et al., 2020), NGRC (Gauthier et al., 2021)), and concurrent time-series LLM-based models (i.e., GPT4TS (Zhou et al., 2023) and Time-LLM (Jin et al., 2023)) as baseline models for performance comparison on proposed LPI4AI dataset.
Experimental Setup. Our experiments are trained with 100 epochs and a batch size of 5, which is adequate to achieve convergence based on our empirical findings. In addition, we utilize a fixed learning rate of $0.0004$ and the Adam optimizer (Kingma & Ba, 2015). A loss function defined by the cumulative absolute error across each time-steps $f_{\text{loss}}(Y,G)=\sum_{n=1}^{\text{pred\_len}}|y_{n}-g_{n}|$ is used, where $Y$ and $G$ represent the sequences of predictions and ground truths, respectively, $y_{n}$ and $g_{n}$ denote the values at the $n$-th time-step, and pred_len is the length of prediction. For all other counterpart methods, we follow their original settings and training configurations to reproduce the results.
@@
Table 1: Quantitative results on LPI4AI test split for hot electron energy predictions (see §3.1 for details). Refer to the metrics section for details of CAE, top-1 and top-5 MAE.
@@
PIC Simulation 2.88 0.20 0.13
LSTM 5.82±0.06 0.35±0.01 0.35±0.01
Autoformer 5.79±0.04 0.35±0.01 0.34±0.01
HoGRC 4.20±0.79 0.25±0.05 0.22±0.02
RCRK 4.31±0.46 0.28±0.04 0.22±0.01
NGRC 4.28±0.68 0.27±0.04 0.23±0.02
GPT4TS 3.34±0.58 0.18±0.05 0.14±0.04
Time-LLM 3.48±0.72 0.18±0.05 0.15±0.05
LPI-LLM (Gemma-2-9B) 2.04±0.21 0.14±0.01 0.12±0.01
LPI-LLM (OLMo-7B) 1.97±0.28 0.14±0.01 0.12±0.01
LPI-LLM (Llama-2-7B) 2.15±0.26 0.14±0.01 0.12±0.01
LPI-LLM (Llama-3-8B) 1.90±0.33 0.14±0.01 0.11±0.01
Performance Comparison. The primary challenge in forecasting for LPI lies in developing a robust yet flexible predictive model. We integrated four wellk-known open-sourced LLMs with our apporach, including Gemma-2 (Team, 2024), OLMo (Groeneveld et al., 2024), Llama-2 (Touvron et al., 2023b), and Llama-3 (AI, 2024). As shown in Table 1, empirical results demonstrate the effectiveness of our method in predicting hot electron energy output in ICF, consistently outperforming all baseline models across all evaluation metrics. Notably, our approach over Llama-3 achieves the best performance and surpasses PIC simulation model (Cao et al., 2022) by 0.98 in terms of CAE. Additionally, compared to classic AI methods, our approach outperforms LSTM (Hochreiter & Schmidhuber, 1997) and Autoformer (Wu et al., 2021) by 3.92 and 3.89, respectively. Our margins in CAE over three other reservoir computing methods, HoGRC (Li et al., 2024), RCRK (Dong et al., 2020), NGRC (Gauthier et al., 2021), are 2.30, 2.41, and 2.38, respectively. Moreover, our method outperforms concurrent LLM-based approaches GPT4TS (Zhou et al., 2023) and Time-LLM (Jin et al., 2023) by 1.44 and 1.58 on CAE, respectively, while maintaining comparable speed. Detailed analysis is supplemented in §S2.
@@
We further present qualitative results in Fig. 5, aligning with our quantitative findings that our method surpasses all comparative baselines in predictive accuracy. Additionally, Fig. 6 illustrates the confidence scores associated with our predictions. The visualization elucidates a clear correlation between predictive error and confidence scores, indicating high confidence corresponding to low errors and conversely. Notably, our approach consistently demonstrates a heightened level of confidence, particularly in forecasting peak values across the sequence, a critical phase in ICF.
Figure 6: Visualization of confidence score and prediction error.
3.2 Diagnostic Experiment
This section ablates LPI-LLM’s systemic design on val split of LPI4AI. All experiments use the Llama 3 8B variant. Appendix §S2 has more experimental results.
Key Component Analysis. We first investigate the two principal modules of LPI-LLM, specifically, Fusion-Specific Prompt and Signal-Digesting Channels. We construct a baseline model with generic dummy prompts which only provide broad, non-specific instructions regarding fusion, and a rudimentary encoder composed of a single linear layer. As shown in Table 2a, the baseline model achieves 3.57 CAE. Upon applying Fusion-Specific Prompt to the baseline, we observe significant improvements for CAE from 3.57 to 2.59. Furthermore, after incorporating Signal-Digesting Channels into the baseline model, we achieve significant gains of 1.56 CAE. Finally, by integrating both core techniques, our LPI-LLM delivers the best performance of 1.19 CAE. These findings affirm that the proposed Fusion-Specific Prompt and Signal-Digesting Channels operate synergistically, and validate the effectiveness of our comprehensive model design.
Table 2: A set of ablative studies evaluated on val split.